OpenStudy (anonymous):

Is there an ISBN that has the form : x - 315266-78-2?

OpenStudy (anonymous):

This question I have right now in my math course, we've had null information about how to do it? could anyone explain to me how to?

OpenStudy (anonymous):

@IrishBoy123 you know anything bout this?

OpenStudy (irishboy123):

ISBN is something to do with books and unique ID numbers and that is everything i know, assuming it is actually correct :-)

OpenStudy (astrophysics):

I don't see how this is a question, ISBN's are on the back of your book, is it asking you to get a book with the certain ISBN maybe?

OpenStudy (anonymous):

@Astrophysics no, you are suppose to show with math calculation if it is a true ISBN or not!

OpenStudy (anonymous):

@perl

OpenStudy (astrophysics):

Oh, I have no idea then, sorry :\

OpenStudy (anonymous):

It has to do with somekind of sequence, and congruity

OpenStudy (anonymous):

we are studying residue class arithmetic or what it is called in english

OpenStudy (perl):

Here is an example ISBN numbers for published books. The ISBN number for our textbook is 0-13-184-868-2. The information is decoded in the first 9 digits. The last digit is for parity check 1 * a1 + 2 * a2 + ... + 9 * a9 = a10 Applying this to our textbook 1 * 0 + 2 * 1 + 3 * 3 + 4 * 1 + 5 * 8 + 6 * 4 + 7 * 8 + 8 * 6 + 9 * 8 = 255 = 2 mod 11 The sum of 9 digits is 255 and it equals to the last digit mod 11.

OpenStudy (anonymous):

⇒ a10 = 4 eftersom 11 | 103 + 40 = 143 \[a _{1}, a _{2}, a _{3}, ..., a _{10}\] the numbers in ISBN \[a _{10}\] is choosen like: \[a _{1}+ 2a _{2}+ 3a _{3}+ ..., 9a _{9}+10a_{10}≡ 0 (mod 11)\] if \[a_{10}=10\] you write X Example ISBN=093603103a10 \[a_{10}\] is choosen like \[1 · 0 + 2 · 9 + 3 · 3 + · · · + 9 · 3 + 10 · a_{10} = 103 + 10a_{10} ≡ 0 (mod 11)\] ⇒ \[a_{10} = 4 \] because \[11 | 103 + 40 = 143\]

OpenStudy (anonymous):

ah wait I gotta write this down and see

OpenStudy (anonymous):

so here I got x+226=2 (mod10)?

OpenStudy (perl):

so we need to check that x - 315266-78-2 can be an isbn number, only if 1*x + 2*3 + 3*1 + 4*5 + 5*2 + 6 * 6 + 7*6 + 8*7 + 9*8 = 2 mod 11

OpenStudy (anonymous):

x+226=2 mod 11

OpenStudy (perl):

check again, I got x + 245 = 2 mod 11

OpenStudy (perl):

1*x + 2*3 + 3*1 + 4*5 + 5*2 + 6 * 6 + 7*6 + 8*7 + 9*8 = x + 245

OpenStudy (anonymous):

yea.. i got the same

OpenStudy (anonymous):

so then should x+245/11 give something with the rest 2..

OpenStudy (perl):

x + 245 = 2 mod 11 x = 2 - 245 mod 11 x = -243 mod 11 x = 10 mod 11

OpenStudy (anonymous):

can you explain the last step?

OpenStudy (perl):

but notice that 10 is not a choice for an isbn digit, since it has to be a digit from 0 to 9 so the answer is, there is no x

OpenStudy (anonymous):

i mean from -243 to 10 mod 11

OpenStudy (anonymous):

yeah but how did you make -243 mod 11 to 10 mod 11?

OpenStudy (perl):

it is easier to use multiple of 11's

OpenStudy (perl):

I am actually using my calculator for negative mod

OpenStudy (perl):

x + 245 = 2 mod 11 with respect to 245 where is the closest multiple of 11 ?

OpenStudy (perl):

22*11 = 242 23*11 = 253

OpenStudy (perl):

245 = 3 mod 11 , so we can substitute (x + 245) mod 11 = (x + 3) mod 11 now solve (x+3) mod 11 = 2 mod 11 x = 2-3 mod 11 x = -1 mod 11

OpenStudy (perl):

when you do negative integers, it is like going backwards with a clock

OpenStudy (perl):

0 mod 11 = 11 mod 11 -1 mod 11 = 10 mod 11 -2 mod 11 = 9 mod 11 -3 mod 11 = 8 mod 11

OpenStudy (perl):